OpenStudy (anonymous):

Which ordered pair is a solution of the equation 2x− y = 9? A)(-4,1) B)(-2,5) C)(5,1) D)(6,-3)

OpenStudy (anonymous):

y=2x-9......any point that lies on this line is a solution im sure you meant to include a list of options...i'll tell you how to do it though....you take those points and plug the x and y values into the equation and see if you get the same thing on both sides...if you do then it is a solution...if not then its not if its just giving a possible solution...pick a random x value and find the corresponding y value...for example when x = 0 2(0) - y = 9 0-y = 9 y=-9 the point (0,-9) is a solution

OpenStudy (anonymous):

@SkylarLee These are all the questions I'm getting from before xD

OpenStudy (anonymous):

I gave options

OpenStudy (anonymous):

Which ordered pair is a solution of the equation 2x − y = 9? (-4,1) (-2,5) (5,1) (6,-3)

OpenStudy (anonymous):

I'm getting (4.5,0), (0, -9)

OpenStudy (anonymous):

I'd go with C

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i got c too

OpenStudy (anonymous):

You can give medals by clicking the "Best Response" button.

OpenStudy (anonymous):

Simplify the expression @Kemist \[3x ^{2}-12x \div4x\]

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

Combine like terms.

OpenStudy (anonymous):

._.

OpenStudy (igreen):

Yep. That leaves you with \(3x^2 - 3x\)

OpenStudy (anonymous):

i got 9x^2-3x

OpenStudy (anonymous):

@iGreen that isnt one of my options :'O

OpenStudy (anonymous):

right wrong yea no ?

OpenStudy (igreen):

What are your options?